OpenStudy (anonymous):

If x=2y-5, and z=16y^3, what is z in terms of x?

OpenStudy (ghazi):

see from above equation y= x+5/2 and we know z=16y^3 therefore z=16(x+5/2)^3\[z=16\frac{ (x+5)^3 }{ (2)^3 }\]

OpenStudy (ghazi):

hope you can break further

OpenStudy (ghazi):

just find y in terms of x and put value of y in the equation Z= 16 Y^3...then use expansion of (a+b)^3

OpenStudy (ghazi):

@mitchelsewbaran can you do it? now it's easier

OpenStudy (anonymous):

hmmm ok